OpenStudy (anonymous):

Solve -2(2x + 5) - 3 = -3(x - 1).

OpenStudy (anonymous):

X=-16

OpenStudy (anonymous):

\[-2(2x+5)-3=-3(x-1)\] by the distribution property \[(-2)(2x)+(-2)(5)-3=(-3)(x)+(-3)(-1)\] \[-4x-10-3=-3x+3\] \[-4x-13=-3x+3\] adding 4x-3 to both sides we get \[-4x-13+4x-3=-3x+3+4x-3\] \[-16=x\]

OpenStudy (anonymous):

−2(2x+5)−3=−3(x−1) Simplify −2(2x+5)−3=−3(x−1) (−2)(2x)+(−2)(5)+−3=(−3)(x)+(−3)(−1)(Distribute) −4x+−10+−3=−3x+3 (−4x)+(−10+−3)=−3x+3(Combine Like Terms) −4x+−13=−3x+3 −4x−13=−3x+3 Add 3x to both sides. −4x−13+3x=−3x+3+3x −x−13=3 Add 13 to both sides. −x−13+13=3+13 −x=16 Divide both sides by -1. −x/−1 = 16 /−1 x=−16